Boulder Furniture has bonds outstonding that mature in 15 years, have a 6 percent coupon, and poy Interest onnually. These bonds have a face value of S1,000 and a current market price of $1075. Whot is the companys oftertox cost of delor if ts tox rote Is 32 percent 0 291 percent O 358 percent O 324 percent 521 percent 553 percent
Answer #1

Coupon = 0.06 * 1000 = 60

Pre tax cost of debt = 5.2645%

Keys to use in a financial calculator: FV 1000, PV -1,075, PMT 60, N 15, CPT I/Y

After tax cost of debt = 0.052645 ( 1 - 0.32)

After tax cost of debt = 0.0358 or 3.58%